# Reasoning or Memorization? Unreliable Results of Reinforcement Learning Due to Data Contamination

### v1

@@ Abstract
The reasoning capabilities of large language models (LLMs) have been a longstanding focus of research within the community. Recent works have further enhanced these capabilities by reinforcement learning (RL), and numerous novel methods have been proposed, claiming to achieve significant improvements with minimal or even no external supervision. Surprisingly, some studies even suggest that random or incorrect reward signals can enhance reasoning performance. However, these breakthroughs are predominantly reported on the Qwen2.5 model family and evaluated on well-known mathematical benchmarks such as MATH-500, AMC, and AIME, while often failing to yield comparable gains on other model families such as Llama, which warrants more in-depth investigation. In this work, our empirical analysis shows that, despite the Qwen 2.5 series attaining superior mathematical-reasoning performance relative to other models, its pre-training on massive web-scale corpora leaves it vulnerable to data contamination in widely used benchmarks (e.g., MATH-500). Consequently, experimental results derived from contaminated benchmarks on the Qwen2.5 series may be unreliable. To obtain trustworthy evaluation signals, we introduce a generator that creates fully synthetic arithmetic problems of arbitrary length and difficulty, yielding clean datasets we call RandomCalculation. Using these leakage-free datasets, we further show that under the RL protocol, only accurate reward signals yield steady improvements that surpass the model’s performance ceiling in mathematical reasoning, whereas noisy or incorrect rewards do not. Thus, we recommend that future studies evaluate on uncontaminated benchmarks and, when feasible, test various model series to ensure trustworthy conclusions about RL and related methods.
@@ 1 Introduction
In recent years, advances in reinforcement learning (RL) techniques have markedly strengthened the reasoning abilities of large language models (LLMs). Flagship systems, including OpenAI’s o1 (OpenAI, 2024b; 2025; a), DeepSeek-R1 (DeepSeek-AI et al., 2025), and QwQ (Qwen Team, 2024; 2025), now match or surpass human-level accuracy on a variety of challenging benchmarks. Among the contenders that excel on these mathematical benchmarks, the open-source Qwen family (Yang et al., 2024a; b; 2025), spanning parameters from 0.5B to 72B and maximally pre-trained on up to 36T high-quality tokens, yields state-of-the-art results in language understanding, mathematics, programming, and preference alignment.
@@
Within this landscape, mathematical reasoning has emerged as a particularly discriminative test bed because it demands precise symbolic manipulation and multi-step logical deduction. Standard suites such as MATH-500 (Hendrycks et al., 2021), AIME (Li et al., 2024), AMC (Li et al., 2024) and Minerva Math (Lewkowycz et al., 2022) require models to parse natural-language problem statements, uncover the latent mathematical structure, and generate exact numeric answers. Recent work has further enhanced these capabilities by reinforcement learning with verifiable rewards (RLVR) (DeepSeek-AI et al., 2025): rule-based reward functions that return 1 when a predicted answer equals ground truth and 0 otherwise. Because the reward is computed analytically, RLVR removes the need for a separate learned reward model, lowering computational cost while providing an analytically precise training signal, especially attractive for domains like mathematics, where solutions are unambiguous.
@@
Although RL nominally depends on accurate reward signals to guide effective training, recent studies (Zuo et al., 2025; Shao et al., 2025) report that even random or partially incorrect rewards can lift Qwen’s performance on standard math benchmarks, whereas the same procedures offer little or no benefit to Llama3.1-8B (Dubey et al., 2024) models. To understand why these seemingly reward-agnostic variants help Qwen but not Llama, we undertake a systematic comparison of the two model families under identical training protocols. We consider two working hypotheses to explain this phenomenon. (i) Data Contamination: Considering Qwen2.5 is pre-trained on massive web-scale corpora from the Internet, including GitHub repositories that store benchmark problems alongside their official solutions. If segments of the evaluation benchmarks leaked into Qwen’s pre-training corpus, spurious rewards could cue the model to retrieve memorized answers rather than acquire new reasoning skills. (ii) Strong Baseline Math Skills: Qwen’s pre-training endows it with higher mathematical capacity than Llama, so even noisy policy-gradient updates appear to help on MATH-500; however, if strong skills are the real driver, the same random-reward trick should still work on a clean benchmark. Distinguishing between these possibilities requires both a leakage audit and a rigorously out-of-distribution (OOD) RLVR evaluation, which we present in Section 4.2 and Section 4.3.
@@
To assess the extent of potential data contamination in popular mathematical benchmarks, we evaluate two indicators: partial-prompt completion rate (can the model reconstruct the tail of a problem?) and partial-prompt answer accuracy (can the model give the correct answer with an incomplete problem?). As Figure 1 illustrates, we observe that Qwen can accurately complete the problem statement and provide the correct answer, whereas Llama does not. Specifically, prompting Qwen2.5-Math-7B with the first 60% of each MATH-500 question, we find that it regenerates the remaining 40% with a 54.60% exact-match rate and still answers 53.6% of these incomplete problems correctly. Llama3.1-8B, in contrast, scores 3.8% and 2.4% on both metrics. Crucially, on the newly released LiveMathBench (version 202505) (Liu et al., 2024), a post-hoc benchmark compiled after the release of the Qwen2.5 model family, Qwen’s completion rate drops sharply to 0.0%, consistent with Llama’s 0.0%. Its partial-prompt answer accuracy also falls to just 2.0%, comparable to Llama’s 1.0%. These results confirm that the earlier gains on MATH-500 may stem from memorized content rather than genuine reasoning. Hence, results derived from MATH-500 and similar datasets for Qwen models should be interpreted with caution.
@@
Building on this evidence, we attribute that data contamination is the chief factor behind the seemingly “magical” success of random-reward or few-shot RLVR variants on Qwen models. To test this claim, we create a fully fresh benchmark (representative examples shown in Figure 2): We use an automatic generator to construct arithmetic expressions of arbitrary length with uniformly random operands and operators, guaranteeing that every instance post-dates the public release of Qwen2.5. A zero-shot evaluation on this benchmark shows no memorization: the accuracy of Qwen2.5 models declines monotonically with the number of computation steps, leaving ample room for improvement. To isolate the effect of reward quality, we next trained Qwen2.5-Math-7B under the standard RLVR protocol on two leakage-free subsets, respectively. The outcome is unambiguous: Correct rewards deliver consistent performance gains, surpassing the model’s performance ceiling; In contrast, random rewards make training highly unstable, yielding no reliable improvement, while inverse rewards rapidly erode the model’s mathematical-reasoning ability. These results rule out the “Strong Baseline Math Skills” explanation and directly implicate “Data Contamination”: once leakage is removed, the prior gains evaporate.
The contributions of our work can be summarized as follows:
We conduct a systematic leakage audit of widely used math benchmarks and demonstrate that Qwen’s sudden performance surge on MATH-500 under spurious-reward RLVR is chiefly attributable to unfair evaluation protocol, data contamination, and memorization rather than its strong math skills.
@@
We design an automatic generator that creates arbitrarily long arithmetic expressions and a corresponding out-of-distribution evaluation protocol. Zero-shot results on this clean suite expose the absence of memorization and place Qwen and Llama on equal footing, enabling fair assessment of mathematical reasoning.
Using the leakage-free dataset, we conduct controlled RLVR experiments and demonstrate that only correct rewards yield stable performance gains, whereas random or inverse rewards provide no benefit, thereby highlighting the central role of reward fidelity. We recommend that future work rely on uncontaminated benchmarks or test various model series to draw trustworthy conclusions about RL-related methods.
@@ 2 Related Works
2.1 Reinforcement learning on Qwen2.5 for mathematical reasoning
A growing body of work investigates how reinforcement learning (RL) can amplify the mathematical-reasoning skills of the open-source Qwen2.5 family. Early studies used explicit verifiable rewards that score an answer as 1 / 0 by exact numerical agreement. Test-time RL (TTRL) (Zuo et al., 2025) applies this signal on-the-fly during inference and yields sizeable gains on MATH-500 and AIME2024. Subsequent efforts pursue extreme data efficiency: a single labeled example can suffice to boost performance (Wang et al., 2025a; see also the confidence-guided RLSC scheme of Li et al., 2025) or even to learn from an unlabeled instance via entropy minimization (Gao et al., 2025). A parallel line replaces external supervision with intrinsic signals derived from the model itself. Entropy-based objectives—RENT (Prabhudesai et al., 2025) and EM-RL (Agarwal et al., 2025)—reward low-entropy output distributions, while other methods such as Intuitor (Zhao et al., 2025b) and online self-training (Shafayat et al., 2025) rely on self-consistency or “self-certainty” as feedback. These approaches report large jumps on Qwen2.5-Math-7B, occasionally matching or surpassing stronger baselines (e.g. GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024)). Other variants explore noisy or partial rewards (Lv et al., 2025) and a label-free RLVR paradigm dubbed Absolute Zero (Zhao et al., 2025a). The generality of these gains, however, has been questioned. Zhang et al. (2025) observe an “underconfidence-to-overconfidence” trajectory in RLIF(Reinforcement Learning from Internal Feedback): accuracy rises early but later collapses as policy entropy shrinks. Shao et al. (2025) further shows that random or incorrect rewards can still improve Qwen2.5 yet fail to transfer to Llama or OLMo (OLMo et al., 2025), suggesting model-specific idiosyncrasies. Together, these findings indicate that while RL can unlock strong mathematical competence in Qwen2.5, its effectiveness hinges on careful reward design, entropy control, and underlying model capacity.
2.2 The Impact of Pretraining Corpora on the Reasoning Ability
The choice of pretraining corpora plays a crucial role in shaping the reasoning abilities of LLMs, particularly in mathematical domains. Several math-specific datasets (Paster et al., 2024; Han et al., 2024; Wang et al., 2024; Allal et al., 2025) have been proposed and have been shown to significantly enhance performance on relevant benchmarks. As highlighted in prior studies, models such as Qwen and Llama, despite having comparable size, exhibit notable differences in mathematical reasoning capabilities. For example, Qwen models tend to benefit more readily from reinforcement learning, whereas Llama models often show limited improvement under similar conditions. Wang et al. (2025b) found that mid-training Llama models on high-quality mathematical corpora substantially improve their performance, both at the base level and after reinforcement learning. These observations suggest that differences in pretraining data may be a key factor influencing a model’s receptiveness to reinforcement learning, and warrant further investigation.
@@ 3.1 Model Selection
Prior research on mathematical reasoning with large language models has focused predominantly on the Qwen-2.5 family (Yang et al., 2024a; b). Accordingly, we center our study on four representative checkpoints from this series: Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-Math-7B, and Qwen2.5-Math-7B-Instruct. For a controlled comparison, we also evaluate Llama3.1-8B and Llama3.1-8B-Instruct (Dubey et al., 2024), which possess comparable parameter counts and thus help isolate model-specific differences in behavior.
3.2 Memorization Capability
We provide the following details on evaluating the model’s memorization capability. Performance is assessed with two metrics: Partial-Prompt Completion Rate and Partial-Prompt Answer Accuracy.
3.2.1 Partial-Prompt Completion Rate
ROUGE Metric.
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) (Lin, 2004) is a family of metrics commonly used to evaluate automatic summarization and text generation systems. ROUGE score measures the overlap between the generated text and one or more reference texts, focusing primarily on recall. The most frequently used variants include:
ROUGE-N: Measures n-gram overlap. ROUGE-1 and ROUGE-2 refer to unigram and bigram recall respectively.
ROUGE-L: Measures the longest common subsequence (LCS) between the generated and reference text, capturing fluency and sentence-level structure.
ROUGE-Lsum: A variant of ROUGE-L adapted for multi-sentence summarization evaluation.
In this work, we utilize average ROUGE-L score to measure the model’s ability to reconstruct the remaining parts of a problem based on partial prefixes, serving as an indicator of the model’s memorization capacity.
Exact Match (EM).
EM is a binary accuracy metric that checks whether the model’s continuation exactly reproduces the reference. For each instance, let denote the model-generated continuation and denote the ground-truth continuation. if , ; otherwise . We first compute the rouge -L score between the and the , then assign
The final EM score is obtained by averaging over all test instances. Because ROUGE-L = 1 implies an exact, character-level match, higher EM values directly indicate a greater proportion of partial-prompts that the model recalls verbatim.
3.2.2 Partial-Prompt Answer Accuracy
Answer-Match Accuracy.
For each question, we supply the model with only a truncated prompt (e.g., the first of the original problem) and allow it to generate an unconstrained continuation. After generation, we check whether the completion contains the ground-truth answer; if so, the instance is scored as correct. Answer-Match Accuracy is defined as the fraction of prompts for which the model’s continuation embeds the correct answer. A high accuracy indicates that the model frequently “recovers” the answer even from a partial problem, which in turn may signal data contamination.
@@ 3.3 RLVR-Based Evaluation
The GRPO Algorithm.
We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as our RLVR algorithm. Formally, for each question , GRPO samples a group of outputs from the old policy and then optimizes the policy model by maximizing the following objective:
where and are hyperparameters, and is the advantage calculated based on the relative rewards of the outputs inside each group only.
Reward Function for RandomCalculation.
The ground-truth answers to our randomly generated arithmetic problems often contain high-precision decimals. When the standard RLVR framework supplies only binary feedback (0 or 1), the model almost never receives positive reinforcement, making training highly unstable and prone to divergence. To address this limitation, we design a continuous reward that ranges from 0 to 1 and jointly penalizes both absolute and relative errors between the model prediction and the reference answer. This denser signal greatly stabilizes reinforcement learning. Let be the model output, be the reference answer, and be a small constant for numerical stability. The reward is computed as:
Reward Design.
Based on the configuration of Shao et al. (2025), we consider the following rewards for our RLVR experiments:
Correct: assigns 1 to a correct answer and 0 otherwise.
@@
Mv-incorrect: retains only majority-voted incorrect labels and assigns 1 when the model output matches an incorrect label, and 0 otherwise.

### v2

@@ Abstract
Reasoning in large language models has long been a central research focus, and recent studies employing reinforcement learning (RL) have introduced diverse methods that yield substantial performance gains with minimal or even no external supervision. Surprisingly, some studies even suggest that random or incorrect reward signals can enhance performance. However, these breakthroughs are predominantly observed for the mathematically strong Qwen2.5 series on benchmarks such as MATH-500, AMC, and AIME, and seldom transfer to models like Llama, which warrants a more in-depth investigation. In this work, our empirical analysis reveals that pre-training on massive web-scale corpora leaves Qwen2.5 susceptible to data contamination in widely used benchmarks. Consequently, conclusions derived from contaminated benchmarks on Qwen2.5 series may be unreliable. To obtain trustworthy evaluation results, we introduce a generator that creates fully clean arithmetic problems of arbitrary length and difficulty, dubbed RandomCalculation. Using this leakage-free dataset, we show that only accurate reward signals yield steady improvements that surpass the base model’s performance boundary in mathematical reasoning, whereas random or incorrect rewards do not. Moreover, we conduct more fine-grained analyses to elucidate the factors underlying the different performance observed on the MATH-500 and RandomCalculation benchmarks. Consequently, we recommend that future studies evaluate models on uncontaminated benchmarks and, when feasible, test various model series to ensure trustworthy conclusions about RL and related methods. Code is available at https://github.com/wumingqi/LLM-Math-Evaluation.
@@ 1 Introduction
In recent years, advances in reinforcement learning (RL) have markedly strengthened the reasoning abilities of large language models (LLMs). Flagship systems, including OpenAI’s o1 (OpenAI, 2024b; 2025; a), DeepSeek-R1 (DeepSeek-AI et al., 2025), and QwQ (Qwen Team, 2024; 2025), already match or exceed human-level accuracy on a variety of challenging benchmarks. Among open-source contenders that excel in these mathematical benchmarks, the Qwen family (Yang et al., 2024a; b; 2025), ranging from 0.5B to 72B and pre-trained on up to 36T high-quality tokens, produces state-of-the-art results in language understanding, mathematics, programming, and preference alignment.
@@
Within this landscape, mathematical reasoning emerges as a particularly discriminative test bed because it demands precise symbolic manipulation and multi-step logical deduction. Standard suites such as MATH-500 (Hendrycks et al., 2021), AIME (Li et al., 2024), AMC (Li et al., 2024), and Minerva Math (Lewkowycz et al., 2022) require models to parse natural-language problem statements, uncover the latent mathematical structure, and generate exact numeric answers. Recent work has further enhanced this capability by reinforcement learning with verifiable rewards (RLVR) (DeepSeek-AI et al., 2025): rule-based reward that returns 1 when the predicted answer equals ground truth and 0 otherwise. Because the reward is computed analytically, RLVR removes the need for a separate learned reward model, lowering computational cost while providing a precise training signal, especially attractive for domains like mathematics, where solutions are unambiguous.
@@
Although RL nominally depends on accurate reward signals to guide training, recent studies (Zuo et al., 2025; Shao et al., 2025) find that even random or incorrect rewards can improve Qwen’s performance on standard math benchmarks, while the same procedures offer little or no benefit to Llama3.1-8B (Dubey et al., 2024). To understand why these seemingly problem-agnostic rewards help Qwen but not Llama, we undertake a systematic comparison of two model families under identical training protocols. We consider two working hypotheses to explain this phenomenon. (i) Data Contamination: Considering Qwen2.5 is pre-trained on massive web-scale corpora from the Internet, including GitHub repositories that store benchmark problems alongside their official solutions. If segments of the evaluation benchmarks leaked into the pre-training corpus, spurious rewards could cue the model to retrieve memorized answers rather than acquire new reasoning skills. (ii) Strong Math Capacity: Qwen’s pre-training endows it with better mathematical capacity than Llama, so even noisy policy-gradient updates appear to help on MATH-500. However, if strong capacity is the real driver, the same spurious rewards should still work on a clean benchmark. Distinguishing between these possibilities requires both a leakage audit and a rigorously out-of-distribution RLVR evaluation.
@@
To assess the extent of potential data contamination in popular mathematical benchmarks, we propose two metrics: partial-prompt completion rate (can the model reconstruct the tail of a problem?) and partial-prompt answer accuracy (can the model give the correct answer with an incomplete problem?). As shown in Fig. 1, Qwen can indeed complete the problem accurately and provide the correct answer, whereas Llama fails. Furthermore, prompting with the first 60% of each MATH-500 problem, Qwen2.5-Math-7B regenerates the remaining 40% with a 54.60% exact-match rate and answers 53.6% of these incomplete problems correctly. In contrast, Llama3.1-8B scores 3.8% and 2.4% on both metrics. Crucially, on the newly released LiveMathBench (version 202505) (Liu et al., 2024), Qwen’s completion rate drops sharply to 0.0%, consistent with Llama’s 0.0%. Its partial-prompt answer accuracy also falls to just 2.0%, comparable to Llama’s 1.0%. These results confirm that Qwen’s pre-training corpus suffers from test data contamination. So, results derived from MATH-500 and similar datasets for Qwen should be interpreted with caution.
@@
Based on this, we attribute that data contamination is the main factor behind the ‘magical’ success of spurious rewards on Qwen. To test this claim, we first create a clean benchmark (i.e., RandomCalculation, example shown in Fig. 2): We use automatic generator to construct arithmetic expressions of arbitrary length with random operands and operators, guaranteeing that every instance post-dates the public release of Qwen. Zero-shot evaluation on this benchmark shows no memorization: the accuracy of Qwen2.5 declines monotonically with the number of computation steps. To isolate the effect of rewards, we next trained Qwen2.5-Math-7B under the standard RLVR protocol on two subsets. The outcome is unambiguous: Correct rewards deliver consistent performance gains, surpassing the model’s performance ceiling. In contrast, random rewards make training highly unstable, yielding no reliable improvement, while inverse rewards rapidly erode the model’s mathematical-reasoning ability. These results rule out the ‘Strong Math Capacity’ hypothesis and directly imply ‘Data Contamination’: once leakage is removed, spurious gains evaporate.
To further test this hypothesis, we measure the similarity between model outputs before and after RL. In MATH-500, the pre- and post-RL responses exhibited substantially higher lexical overlap than in RandomCalculation, indicating that Qwen inadvertently retrieves its memory and answers during RL with spurious rewards. A more detailed token-level analysis also supports this: the token-level KL divergence between the pre- and post-RL models is significantly lower for MATH-500. These results strengthen our hypothesis that data contamination leads to successful RL through spurious rewards on Qwen series. Based on these findings, we recommend that future work should test on uncontaminated benchmarks or more diverse model series to draw trustworthy conclusions about RL-related methods.
Contributions of our work can be summarized as follows:
We conduct a systematic leakage audit of math benchmarks with two novel metrics and demonstrate that Qwen suffers from data contamination on public benchmarks.
@@
We propose an automatic generator that creates arbitrarily long arithmetic expressions. Zero-shot evaluation on this dataset exposes the absence of memorization, enabling fair assessment of RL methods.
Using this clean dataset, we conduct RL experiments and demonstrate that only correct reward yields stable improvement, whereas spurious rewards provide no benefit.
We reveal that spurious rewards solely enable the retrieval of memory from pre-training, leading to spurious performance improvement on MATH-500.
@@ 2 Related Works
2.1 RL on Qwen2.5 for Mathematical Reasoning
A growing body of work investigates how reinforcement learning (RL) can amplify the mathematical-reasoning capacity of the open-source Qwen2.5 family. Early studies use verifiable rewards that score an answer as 1 / 0 by exact numerical agreement. Test-time RL (Zuo et al., 2025) applies this signal on-the-fly during inference and yields sizeable gains on MATH-500 and AIME2024. Subsequent efforts pursue extreme data efficiency: few labeled (Wang et al., 2025a; Li et al., 2025) or even unlabeled examples (Gao et al., 2025; Zhao et al., 2025a) can suffice to boost performance. A parallel research replaces external supervision with intrinsic signals derived from the model itself: Prabhudesai et al. (2025); Agarwal et al. (2025) reward low-entropy output distributions, while Zhao et al. (2025b); Shafayat et al. (2025) rely on self-consistency or self-certainty as feedback. These approaches report large jumps on Qwen2.5-Math-7B, occasionally matching or surpassing stronger supervised baseline. Other variants explore noisy (Lv et al., 2025) or even random rewards (Shao et al., 2025) for Qwen2.5. However, these methods fails to transfer to Llama or OLMo (OLMo et al., 2025), suggesting model-specific idiosyncrasies.
2.2 Factors Influencing Performance on Math
The choice of pretraining corpora plays a crucial role in shaping the reasoning abilities of LLMs, particularly in mathematical domains. Several math-specific datasets (Paster et al., 2024; Han et al., 2024; Wang et al., 2024; Allal et al., 2025; Stojanovski et al., 2025) have been proposed and shown to significantly enhance performance on relevant benchmarks. Further, Wang et al. (2025b) finds that mid-training Llama models on high-quality mathematical corpora substantially improve their capacity, both at the base level and after reinforcement learning. However, evaluation of math capability can be misleading due to potential test data contamination. Xu et al. (2024) observed that certain widely-used benchmarks may be partially included in the pretraining corpus of LLMs, including early versions of Qwen. On the other hand, Liu et al. (2025) found that omitting dialogue-style prompting can lead to improved mathematical reasoning. These observations underscore the presence of multiple confounding factors in evaluating model performance, motivating the need for rigorous and systematic analysis.
@@ 3.1 Model Selection
Prior researches on mathematical reasoning with LLMs focus predominantly on the Qwen-2.5 (Yang et al., 2024a; b). Accordingly, we center our study on four representative checkpoints from this series: Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-Math-7B, and Qwen2.5-Math-7B-Instruct. For a controlled comparison, we also evaluate Llama3.1-8B and Llama3.1-8B-Instruct (Dubey et al., 2024), which possess comparable parameter counts and thus help isolate model-specific differences in behavior.
3.2 Evaluation of Memorization Capability
We assess the model’s memorization of benchmark data, i.e., data contamination, using two metrics: Partial-Prompt Completion Rate and Partial-Prompt Answer Accuracy.
Specifically, we prompt the model to complete the remaining parts of a problem based on partial prefixes. To evaluate its performance, we use the Partial-Prompt Completion Rate measured by ROUGE-L (Lin, 2004), which calculates the overlap of the longest common subsequence between the generated and reference text, capturing fluency and sentence-level structure. Additionally, we utilize Exact Match (EM) accuracy, which checks if the model’s output exactly matches the reference. The final EM score is the average across all instances. Higher EM indicates a greater proportion of partial-prompts that model can recall exactly.
Besides, for each question, we supply the model with only a truncated prompt (e.g., the first of the original problem) and allow it to generate an unconstrained continuation. After generation, we check whether the completion contains the ground-truth answer; if so, the instance is scored as correct. Partial-Prompt Answer Accuracy is defined as the fraction of prompts for which the model’s continuation embeds the correct answer. A high accuracy indicates that the model frequently ‘recovers’ the answer even from a partial problem, which in turn may signal data contamination.
@@ 3.3 RLVR-Based Evaluation
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is adopted as our RLVR algorithm. Formally, for each question , GRPO samples a group of outputs from the old policy and then optimizes the policy model by maximizing the following objective:
where , and are hyperparameters, and is the advantage calculated based on the relative rewards of the outputs inside each group only.
Spurious Reward.
Following Shao et al. (2025), we consider following spurious reward types for RLVR:
@@
Mv-incorrect: uses the majority-voted incorrect labels from the model, assigning a reward of 1 when the model output matches an incorrect label, and 0 otherwise.